You time an ant’s movement as it walks along a meter stick. You record the following data: {(0 s, 0.20 m), (1 s, 0.26 m), (2 s, 0.32 m), (3 s, 0.38 m), (4 s, 0.44 m)}. Determine the ant’s velocity.


1
Expert's answer
2021-10-28T15:21:40-0400

Let's plot the Position versus Time graph:



As we can see, the Position vs. Time graph is a straight line. The slope of the Position vs. Time graph represents the velocity of the ant. Let's find the slope of the graph:


"v_{ant}=Slope=\dfrac{\Delta y}{\Delta x}=\dfrac{0.44\ m-0.2\ m}{4\ s-0\ s}=0.06\ \dfrac{m}{s}."

Answer:

"v_{ant}=0.06\ \dfrac{m}{s}."